OpenStudy (anonymous):

k=3w+2 In the equation above, if w is increased by 2 by how much does k increase?

OpenStudy (anonymous):

since w is increased by 2, then \(\large k=3(w+2)+2=3w+6+2=(3w+2)+6 \) so (3w+2)+6=k+6 \)

OpenStudy (anonymous):

thank you !